If Negan Corp. common stock is valued at $40 per share, dividends are paid quarterly and expected to
grow quarterly by 1.2% forever, and the next dividend is due in 3 months and expected to be $3, then
what is the expected annual return on Negan Corp. stock?
1
Expert's answer
2020-04-13T19:31:39-0400


find current profitability

"r+q=\frac{d}{P}=r+0.012=\frac{3}{40}"

r=0.063, 6,3%


find annual dividends:

  1. "D=D1\times(1+q)^4=3\times(1+0.012)^4=3.146"


find the annual yield:

"r=\frac{d}{P}=\frac{3.146}{40}=0.07866" , 7.87%